# A method to benchmark high-dimensional process drift detection

### v1

@@ Abstract.
Process curves are multi-variate finite time series data coming from manufacturing processes. This paper studies machine learning methods for drifts of process curves. A theoretic framework to synthetically generate process curves in a controlled way is introduced in order to benchmark machine learning algorithms for process drift detection. A evaluation score, called the temporal area under the curve, is introduced, which allows to quantify how well machine learning models unveil curves belonging to drift segments. Finally, a benchmark study comparing popular machine learning approaches on synthetic data generated with the introduced framework shown.
@@ 1. Introduction
Manufacturing lines are typically decomposed into a sequential arrangement of individual steps, called processes, each utilizing one or more manufacturing techniques, such as casting, forming, separating, joining, or coating, to develop the component in accordance with its final specifications. Modern and encompassing sensor technology allows to monitor process behavior precisely, typically by measuring key performance indicators, like force, pressure, or temperature, over time. In modern IoT-enabled manufacturing lines, persisting these process curves has become the general case, allowing to monitor not only the process behavior when working on a single component, but also globally over multiple sequential executions [32]. Malign events like anomalous component batches, tool wear, or wrong calibrations can affect the line performance badly. Such events proceed subtle and often lead to a slow deformation of the resulting process curve iteration by iteration. Detecting such process drifts, that is time periods where machine behavior changes, is key to keep unplanned downtimes and costly bad parts at bay [19].
Specifically in high-volume production settings, where processes yield curves with multiple variables at a small rate and with short cycle times between subsequent executions, detecting process drifts is challenging due to the vast amount of data that is generated in short time. Thus, these settings have been an ideal application for machine learning methods [9, 15, 19, 30, 31, 2, 21]. Process drifts should not confused with concept or data drifts, where the goal is typically to analyse the declining performance of a trained machine learning model when new data starts to differ from the train data [22]. Thus, methods to detect concept drifts often have access to train data and trained machine learning models that can be used. In our setting, however, we are interested in machine learning models that detect the drift itself. Generally speaking, process curve datasets are datasets holding finitely many multi-variate finite time series. Although process curves are time series by definition, many analysis techniques to analyse time series data do not apply because many methods, like ARMA [7] methods, assume infinite time series, often with strong assumptions on their stationarity.
When working with high-dimensional data, multivariate functional data analysis [16, 24] is the statistical workhorse. Deep learning techniques, particularly dimensionality reduction methods like autoencoder [25, 14], become increasingly popular [20]. A popular approach in practice is to learn an low-dimensional embedding of the high-dimensional input and then analyse the learned latent variables with classic machine learning models. There exist some architectures that also can directly consume temporal context, like the index of the process iteration, as an auxiliary variable to learn causal meaningful latent variables, like the iVAE [12] or the independent component analysis [11]. In [13], a derivate of an variational autoencoder has been developed which particularly deals with process drifts. In [31], Bayesian autoencoders where used to detect drifts in industrial environments. Moreover, more general neural network based systems, like for casting [15] or milling [9] have been developed to analyse process curves. Often, also expert knowledge can be utilized, like by extracting deterministic features which then are further analyzed by machine learning methods, like in [2] for process curve analysis of wire bounds. Once the dimensionality has been reduced, either by hand-crafted feature extraction of representation learning, classic methods for drift detection apply, like a sliding KS-test [23]. There are also methods that work in high-dimensional settings for tabular data, like hellinger-distance based techniques [5].
A common challenge when applying machine learning methods on applications from manufacturing is data imbalance as interesting events like scrap parts or drifts are rare. In [27], a deep generative model was used to synthetically generate process curves for casting processes, like to increase observations from the minority class.
@@
Overall, the analysis of process curves with machine learning techniques has been studied in depth for lots of applications from manufacturing. However, some inherent challenges remain. First, whether a model that works well in one application also does in another is by far not guaranteed. Thus, releasing data to the public and allowing to benchmark multiple models on the same data is important to make machine learning research more transparent for process curve analysis. Most papers mentioned above, no data has been released to the public, making it impossible to test other architectures. Often, this is due to privacy issues and fear of leaking information to competitors. Public real-world datasets to study process drifts are rare. For the special situation of milling, two process curve datasets [1, 28] exist that hold ground truth quantifying the health of the tool used. A comprehensive overview of degradation data sets for tasks with for tasks within prognostics and health management can be found in [18]. For causal discovery, a few approaches exist that realistically generated synthetic data, like for time-series [4] and tabular production data [8], however, the data generated is often identically and independently distributed (iid) and hence inappropriate for drift detection. The second challenge is how models should be evaluated. In most drift detection settings, data is imbalanced, meaning that there are far less drifts than normal data points. In [13], the false alarm rate and accuracy have been used to evaluate the model performance. A commonly used metric to measure the statistical performance of a binary classifier, like an anomaly detector, is the area under the ROC curve - short AUC. However, the usage of the AUC is typically only applicable in settings where data is assumed to be iid, which is not the case in drift detection.
In this work, we address the issue of benchmarking machine learning models for process drift detection. We present a simple, flexible and effective theoretic framework to generate synthetic process curves including drifts with a validated ground truth (Section 2 and Section 3). Its possible to feed data of real process curves making the data realistic. Moreover, we construct a metric called temporal area under the curve (TAUC) in Section 4, which aims to take the temporal context of a detection into account. Finally, we present a short benchmarks study in Section 5 as a proof of concept for the effectiveness of our score in measuring the predictive power of drift detectors. Our work is based on preliminary results of the first author [29]. In this work, we provide additionally insights into the introduced score, introduce a variant called soft TAUC and compare it in depth with existing scores. Moreover, we substantially generalize the data synthetization framework, for instance by allowing higher-order derivatives, and we generate more sophisticated datasets for the benchmark study. We also release the code that helps generating process curves to benchmark drift detectors, which is freely available under https://github.com/edgarWolf/driftbench. Its optimization back-end is implemented in Jax [3] allowing a fast GPU-based curve generation.
2. Statistical framework to analyse process drifts
A process curve is a finite time-series $C=(Y(x))_{x\in I}$ with $Y(x)\in\mathbb{R}^{c}$ and $I\subset\mathbb{R}$ a finite set, where $Y:\mathbb{R}\to\mathbb{R}^{c}$ represent physical properties of the process to be measured and $x$ an independent variable, often the time (see Remark 2.1 for concrete examples). We call $c\in\mathbb{N}$ the dimension of the process curve. Whenever a manufacturing process finishes its work on a component, a process curve is yielded. Thus, when the same process is executed on multiple parts sequentially, a sequence $C_{1},\ldots,C_{T}$ is obtained where each $C_{t}$ arises under slightly different physical conditions, i.e., $C_{t}=(Y_{t}(x))_{x\in I_{t}}$. The different conditions can be due to different properties of the components or due to degradation and wareout effects of tools within the machine running the process. Also, the elements in $I_{t}$ can vary from execution to execution, for instance due to different offsets.
Remark 2.1.
In staking processes $Y$ is a measured force and $x$ the walked path of the press. In pneumatic test stations, $Y$ might be a pressure where $x$ might be time. In bolt fastening processes, $Y$ holds torque and angle and $x$ is the time [21].
As new components are assembled each time the process is executed, the effect of a part only affects a given curve, whereas wareout of the tool, which is used in every process, affects all curves. Let $C_{1},\ldots,C_{T}$ be process curves from $T$-many executions and $Y_{1},\ldots,Y_{T}$ the corresponding physical relations. We write $[T]$ for the set $\{1,\ldots,T\}$. In our approach, we assume that there exists a function $f:\mathbb{R}^{k}\times\mathbb{R}\to\mathbb{R}^{c}$ as well as a function $w:[T]\to\mathbb{R}^{k}$, such that for all $t\in[T]$ and $x\in I_{t}$:
$f(w(t),x+{\epsilon}_{x})+{\epsilon}_{y}=Y_{t}(x)$
where the function $f$ is a proxy for the physics underneath the process and ${\epsilon}_{x}$ and ${\epsilon}_{y}$ denote white noise relating to measurement inaccuracies that is independent of $t$. The vector $w(t)$ represents environmental properties of the $t$-th execution, and some of its coordinates correspond to component properties, some to properties of the machine.
Assuming no tool degradation but only component variance, we could assume that $w(t)$ is sampled in each process from a fixed but unknown distribution on $\mathbb{R}^{k}$, like $w(t)\sim\mathcal{N}_{\mu,\sigma}$ with fixed $\mu\in\mathbb{R}^{k}$ and $\sigma\in\mathbb{R}^{k\times k}$ for all $t\in[T]$. Tool degradation, however, affects the curve substantially, often by letting certain support points of the curve move. For instance, in a staking process, the position and value of the maximal force, i.e., where the first derivative is zero, starts shifting (see Figure 1). Often, these support points can be formulated in terms of derivatives of $f$. More formally, let $\partial^{i}_{x}$ be the $(i)$-th derivative of $f$ according to the second argument and let for $t\in[T]$, $x^{i}(t),y^{i}(t)\in\mathbb{R}^{n_{i}}$ be such support points of the $t$-curve $C_{t}$, i.e.:
(2.1) $\partial_{x}^{i}f(w(t),x^{i}_{j}(t))=y^{i}_{j}(t)$
for all $j\in[n_{i}]$. Particularly, the properties $w(t)$ deform in a way such that the derivatives of $f(w(t),\cdot)$ satisfies all support points. As mentioned before, if underlying physical properties of the process or components change, these support points start to change their position. More formally, at the $t$-th process execution, the support points are sampled according to a distribution, whose statistical parameter depend on $t$, for instance, for a particular support point $j\in[n_{i}]$ of $y^{i}$, the following temporal change could happen, i.e. $y^{i}_{j}(t)\sim\mathcal{N}_{\mu^{i}(t),\sigma}$, with
$\mu^{i}(t)=\begin{cases}b,&\textnormal{ if }t<t_{0}\\ d\cdot\frac{t-t_{0}}{t_{1}-t_{0}}+b,&\textnormal{ if }t_{0}\leq t\leq t_{1}\\ d,&\textnormal{ if }t_{1}<t\end{cases}$
with $b\neq d$. That is, the mean of the support point drifts linearly from $b$ to $d$ between $t_{0}$ and $t_{1}$ (see also Figure 1).
@@ 3. Data generation
In this section, we explain the synthetization method of process curves. Here, we neither focus on how the $w(t)$ behave in latent space, nor on how $f$ is formulated exactly. Instead, our key idea is to model the behavior of support points over time and seek for parameters $w(t)$ using non-linear optimization satisfying the support point conditions in (2.1). That way, we obtain a generic and controllable representation of the process curves that is capable of modelling physical transformations over time by letting the physically motivated support points drift over time. Without restricting generality and to keep notation simple, we will assume for the remainder that $c=1$. The multivariate case is a straight-forward application of our approach by modeling each variable in $Y$ individually (see Remark 3.1). For this, let $f:\mathbb{R}^{k}\times\mathbb{R}\to\mathbb{R}$ be an $l$-times differentiable map and let for each $i\in[l]$, $x^{i}\in\mathbb{R}^{n_{i}}$ and $y^{i}\in\mathbb{R}^{n_{i}}$ be two $n_{i}$-dimensional vectors. We want to find $w^{*}$ such that
(3.1) $\partial^{i}_{x}f(w^{*},x^{i}_{j})\approx y^{i}_{j}\quad\forall j\in[n_{i}]\quad\forall i\in[l].$
Particularly, $x^{i}$ and $y^{i}$ can be considered as support points surpassed by the graph of the map $f(w^{*},\cdot):\mathbb{R}\to\mathbb{R}$ (see Figure 2).
@@
Fig. 2. Visualization of the data synthetization given a function $f(w,x)=\sum_{i=0}^{5}w_{i}\cdot x^{i}$. Left figure shows $f(w,\cdot)$ solved for concrete $x^{i},y^{i}$ (red points). Right figure shows sequence $f(w_{1},\cdot),\ldots,f(w_{100},\cdot)$ where gaussian noise was added on on coordinate in $y^{1}(t)$ (corresponding coordinate $x^{1}(t)$ is marked with a dashed line).
Assuming that $f\in\mathcal{C}^{l+2}(\mathbb{R}^{k}\times\mathbb{R})$, i.e. $f$ is $l+2$-times differentiable, and given $x^{i}\in\mathbb{R}^{n_{i}}$ and $y^{i}\in\mathbb{R}^{n_{i}}$ for each $i\in[l]$, we can solve (3.1) using second-order quasi-Newton methods [6, Chapter 3] for the objective function
(3.2) $\operatorname*{arg\,min}_{w\in\mathbb{R}^{k}}\sum_{i=1}^{l}\sum_{j=1}^{n_{i}}D_{i}\cdot\|\partial^{i}_{x}f(w,x^{i}_{j})-y^{i}_{j}\|^{2}$
where $D_{1},\ldots,D_{l}$ are constants to account for the different value ranges of the functions $\partial^{i}_{x}f$. Now, let $x^{i}(1),\ldots,x^{i}(T)\in\mathbb{R}^{n_{i}}$ and $y^{i}(1),\ldots,y^{i}(T)\in\mathbb{R}^{n_{i}}$ be sequences of $T$-many $n_{i}$-dimensional vectors respectively. Solving the optimization problem (3.2) for the each $t\in[T]$, i.e. for the $x^{1}(t),y^{1}(t),\ldots,x^{l}(t),y^{l}(t)$, we get a sequence of parameters $w_{1},\ldots,w_{T}\in\mathbb{R}^{k}$ and a sequence of maps $f(w_{1},\cdot),\ldots,f(w_{T},\cdot)$ from $\mathbb{R}$ to $\mathbb{R}$.
Fixing now a finite and not necessarily equidistant set of points of interests $I\subset\mathbb{R}$ and setting $C_{i}=f(w_{i},I)\in\mathbb{R}^{|I|}$, we obtain a sequence of process curves $C_{1},\ldots,C_{T}$. In Appendix A, we showcase in depth an example where $f$ is a polynomial.
@@
Of course, our theoretic framework extends naturally to multi-variate time series data, where each signal has either its own or multiple signals share same latent information.
Remark 3.2 (Feeding real-data).
Clearly, the support points $x^{1}(t),y^{1}(t),\ldots,x^{l}(t),y^{l}(t)$ can come from a real process curve dataset. Choosing a universal function approximator $f$ like a feed-forward neural network [10] or a Kolmogorov-Arnold network [17].

### v2

@@ Abstract.
Process curves are multivariate finite time series data coming from manufacturing processes. This paper studies machine learning that detect drifts in process curve datasets. A theoretic framework to synthetically generate process curves in a controlled way is introduced in order to benchmark machine learning algorithms for process drift detection. An evaluation score, called the temporal area under the curve, is introduced, which allows to quantify how well machine learning models unveil curves belonging to drift segments. Finally, a benchmark study comparing popular machine learning approaches on synthetic data generated with the introduced framework is presented that shows that existing algorithms often struggle with datasets containing multiple drift segments.
@@ 1. Introduction
Manufacturing lines typically consist of processes arranged sequentially, each using techniques like casting, forming, or joining to shape components to their final specifications. Advanced sensor technology enables precise monitoring of key performance indicators, like force, pressure, or temperature, over time. IoT-enabled systems now commonly store the data obtained, called process curves, facilitating analysis across both single components and entire production sequences [1]. Issues like anomalous batches, tool wear, or miscalibrations can degrade performance, often subtly, by causing gradual shifts in process curves. Thus, detecting process drifts is key to keep unplanned downtimes and scrap parts at bay. In high-volume production, this is particularly challenging due to the rapid data generation and complexity of multi-variable curves and hence these settings have been an ideal application for machine learning methods [2, 3, 4, 5, 6, 7, 8]. Although process curves are multivariate time-series, process drift detection should not be confused with drift detection in time series [9] or drifts in profile data [10] (see also Figure 1)). Typically, statistical drift detection methods from time series analysis are not direct applicable, not alone because process curves are high-dimensional objects, but also because of high autocorrelation among their sample axis. As a consequence, deep learning techniques, most prominently dimensionality reduction methods like autoencoders [11, 12, 13, 14, 15], have become increasingly popular [16] as they allow to first learn a low-dimensional representation of the high-dimensional input and then analyse the learned latent variables with classic statistical tools, like sliding Kolmogorov-Smirnov tests [17], Hellinger-distance based techniques [18], or by facilitating the Maximum Mean Discrepancy [19].
Time series dataProfile dataProcess curve dataProcess executionsRefer to captionRefer to captionRefer to captionRefer to captiondriftRefer to captiondriftRefer to captionRefer to captionRefer to captionRefer to captiondrift
Fig. 1. Overview of different kinds of time series data from manufacturing processes and drifts within.
For many machine learning applications relevant for manufacturing, established ways and datasets to benchmark the performance of algorithms exist, like for causal discovery in quality data [20], anomaly detection in images from optical inspections [21], or reinforcement learning in continuous control tasks [22]. However, for process drift detection, such a framework is yet missing to the best of our knowledge. This may be due to the following two reasons: The lack of both, publicly available datasets and a suitable evaluation metric. Beside a few publicly released datasets [23, 24, 25], most of the existing work does not release any data to the public, making it impossible to test other detectors on the same dataset. Often, this is due to privacy issues and fear of leaking information to competitors. In addition, as datasets for process drifts are inherently non identically and identically distributed (iid), any sort of test and train splits introduced significant biases making evaluation of algorithms hard if only one variant of a dataset is available Finally, process drift detection is by definition an unsupervised learning task, but to benchmark detectors, a ground truth is required, labeling precisely when a drift starts and when it ends. This seamlessly leads to the second challenge, namely the missing evaluation metric. As in any machine learning task, the metric depends on the precise application and a trustworthy ground truth. A commonly used metric used in research and practice to measure the statistical performance of a binary classifier, often independent of the application, is the area under the ROC curve [26] - short AUC. However, the usage of the AUC is typically only applicable in settings where data is assumed to be iid, unlike in drift detection. In our work, we want to exactly address these issues by introducing a benchmarking framework for researchers, allowing them to reliably validate their process drift detection algorithms. At a high level, our main contributions are:
  • We present a simple, yet flexible and effective theoretic framework to generate synthetic process curve datasets including drifts with a validated ground truth (Section 2 and Section 3) which also allows feeding of curves from real processes.
  • We introduce an evaluation metric called temporal area under the curve (TAUC) in Section 4, which aims to take the temporal context of a detection into account.
  • We conduct a short benchmark study in Section 5 as a proof of concept for the effectiveness of both, our TAUC metric and our proposed data generation method to measure the predictive power of drift detectors.
Our work is based on preliminary results of the first author [27]. In this work, we provide additionally insights into the introduced metric, introduce a variant called soft TAUC and compare it in depth with existing metrics. Moreover, we substantially generalize the data synthetization framework, for instance by allowing higher-order derivatives, and we generate more sophisticated datasets for the benchmark study. We also release the code that helps to generate process curves to benchmark drift detectors, which is freely available under https://github.com/edgarWolf/driftbench. Its optimization back-end is implemented in Jax [28] allowing a fast GPU-based generation of process curves.
2. Statistical framework to model process drifts
In this section, we formalize what we consider as process curves and drifts within. Generally speaking, process curve datasets are datasets consisting of finitely many multivariate time series each having finitely many steps.
@@ 2. Statistical framework to model process drifts
We formally model a process curve as a finite time-series $(Y(x))_{x\in I}$ with $Y(x)\in\mathbb{R}^{c}$, $I\subset\mathbb{R}$ a finite set, and where $Y:\mathbb{R}\to\mathbb{R}^{c}$ represent physical properties of the process to be measured and $x$ an independent variable, often the time. In staking processes, for instance, $Y$ is the measured force and $x$ the walked path of the press (compare also [20, Figure 9]). Another example are pneumatic test stations, where $Y$ might be a pressure measured over time $x$. In bolt fastening processes, $Y$ represents the torque measured over the angle $x$ [8]. We call the number of variables $c\in\mathbb{N}$ in the curve the dimension of the process curve and write $[T]$ for the set $\{1,\ldots,T\}$ and often refer to it as temporal axis.
Whenever a manufacturing process finishes its work on a component, a process curve is yielded. Thus, when the same process is executed on multiple times sequentially, a long sequence $C_{1},C_{2},\ldots,C_{t},\ldots,C_{T}$ with $T\in\mathbb{N}$ of process curves is obtained where each $C_{t}$ arises under slightly different physical conditions $Y_{1},\ldots,Y_{T}$ , i.e., $C_{t}=(Y_{t}(x+{\epsilon}_{x})+{\epsilon}_{y})_{x\in I_{t}}$, where ${\epsilon}_{x}$ and ${\epsilon}_{y}$ represents measurement noise or inaccuracies. In theory, also the sets $I_{t}$ can vary for each $t\in[T]$, for instance due to different offsets. Wearout or tool degradation affects the process curves gradually and to model their deformation along the execution axis, we assume that there exists functions $f:\mathbb{R}^{k}\times\mathbb{R}\to\mathbb{R}^{c}$ and $w:[T]\to\mathbb{R}^{k}$ such that for all $t\in[T]$ and $x\in I_{t}$:
(2.1) $f(w(t),x)=Y_{t}(x)$
where the function $f$ is a proxy for the physics underneath the process. The vector $w(t)\in\mathbb{R}^{k}$ represents environmental properties of the $t$-th execution, and some of its coordinates correspond to component properties, some to properties of the machine. Without restricting generality and to keep notation simple, we will assume for the remainder that $c=1$, as the multivariate case is a straight-forward application of our approach by modeling each variable in $Y$ individually (see also Remark 3.1).
Assuming only component variance and no tool degradation, we could assume that $w(t)$ is sampled in each process execution from a fixed but unknown distribution on $\mathbb{R}^{k}$, like $w(t)\sim\mathcal{N}_{\mu,\sigma}$ with fixed $\mu\in\mathbb{R}^{k}$ and $\sigma\in\mathbb{R}^{k\times k}$ for all $t\in[T]$. As mentioned, tool degradation, in contrast, affects the process from execution to execution, i.e., the parameters of the distribution shift over time leading to a deformation of the observed process curve. Such process drifts should not be confused with concept drifts, where the goal is typically to analyse the declining performance of a trained machine learning model when new data starts to differ from the train data [29]. Moreover, detecting drifts in process curves is different to detecting drift in profile data [10], where one typically is interested in drifts among the curves yielded by a single execution, not in drifts over multiple executions. A similar application is the identification of drifts within profile data, where typically one process execution yields a sequence of process curves of fixed size, like in spectroscopy when one curve is some intensity over time which is measured for different wavelengths [30]. One way to model process drifts is to model the evolution of the latent parameters $w(t)$, like using a dynamical system. For instance, in control theory [31], $w(t)$ is considered as latent state of a system which evolves over the executions $t$ and one observes a multivariate output $Y(t)\in\mathbb{R}^{|I_{t}|}$ with $Y(t)=Y_{t}(I_{t})$. Introducing a control vector $u(t)\in\mathbb{R}^{p}$, $w(t)$ can be considered as state variable $w(t)$ of the system that evolves over time and is influenced by a control vector $u(t)\in\mathbb{R}^{p}$ such that $\partial_{t}w(t)=h(w(t),u(t),t)$ and $Y(t)=f(w(t))$ holds for all $t\in\mathbb{N}$. Here, however, one has to precisely model how the state $w$ changes over executions and how it is affected by interventions $u$ and has to solve challenging non-linear differential equations. However, as we will argue, the degradation of the curve can be described directly in curve space in many scenarios. Thus, we directly model the transformation of the process curves in curve space by letting certain support points of the curve move in a controlled way:
Definition 2.1 (Support points).
Let $f:\mathbb{R}^{k}\times\mathbb{R}\to\mathbb{R}$ be an $i$-times differentiable function, $i\in\mathbb{N}$, and $\partial_{x}^{i}f$ be the $i$-th derivative of $f$ according to the second argument. Let $x,y\in\mathbb{R}^{n}$, then $(x,y)$ is a support point of $i$-th order for $f$ at $w\in\mathbb{R}^{k}$ if $\partial^{i}_{x}f(w,x_{j})=y_{j}$ for all $j\in[n]$.
Support points can be considered as points surpassed by the graph of $f(w,\cdot):\mathbb{R}\to\mathbb{R}$ (see visualization on the left in Figure 3). Typically, such support points are physically motivated and if latent properties of the process change, certain support points change their position in curve space. For instance, in a staking process, the position $x(t)$ and value $y(t)$ of the maximal force, i.e., where the first derivative is zero, starts shifting (see Figure 2). That is, we can describe this behavior by modelling the support points $(x(t),y(t))$ and $(x(t),0)$ of first and second order respectively, i.e. $f(w(t),x(t))=y(t)$ and $\partial_{x}^{1}f(w(t),x(t))=0$. We formalize in Section 3 how we can use this to generate process curves and drifts synthetically.
@@ 3. Data generation
Let $f:\mathbb{R}^{k}\times\mathbb{R}\to\mathbb{R}$ be as in Section 2 a proxy for the physical relations for given manufacturing process. In this section, we build our synthetization framwork upon the setup introduced in Section 2. Here, we neither focus on how $w(t)$ behaves in latent space, nor on how $f$ is formulated exactly. Instead of modeling the evolution of $w(t)$ with a dynamic system, our idea is to model the behavior of support points over process executions in curve space and to seek for parameters $w(t)$ using non-linear optimization satisfying the support point conditions from Definition 2.1. For the remainder of this section, we explain how $w(t)$ can be computed given the support points. Thus, assume we have for each process execution $t\in[T]$ support points $(x^{1}(t),y^{1}(t)),\ldots(x^{l}(t),y^{l}(t))$ with with $x^{i}(t),y^{i}(t)\in\mathbb{R}^{n_{i}}$, that is,
(3.1) $\partial_{x}^{i}f(w(t),x^{i}_{j}(t))=y^{i}_{j}(t)\quad\forall j\in[n_{i}].$
@@
Fig. 3. Visualization of the data synthetization given a function $f(w,x)=\sum_{i=0}^{5}w_{i}\cdot x^{i}$. Left figure shows $f(w,\cdot)$ solved for concrete $x^{i},y^{i}$ (red points). Right figure shows sequence $f(w_{1},\cdot),\ldots,f(w_{100},\cdot)$ where gaussian noise was added on one coordinate in $y^{1}(t)$ (corresponding coordinate $x^{1}(t)$ is marked with a dashed line).
Instead of modelling $w(t)$ explicitly, we compute $w(t)$ implicitly such that (3.1) is satisfied. For instance, if $f$ is $l+2$-times differentiable in its second argument and if $\partial_{w}^{2}\partial_{x}^{i}f$ exists, we can solve (3.1) individually for all $t\in[T]$ using second-order quasi-Newton methods [32, Chapter 3] for the objective function
(3.2) $w(t)=\operatorname*{arg\,min}_{w\in\mathbb{R}^{k}}\sum_{i=1}^{l}\sum_{j=1}^{n_{i}}D_{i}\cdot\left(\partial^{i}_{x}f(w,x^{i}_{j}(t))-y^{i}_{j}(t)\right)^{2}$
where $D_{1},\ldots,D_{l}$ are constants to account for the different value ranges of the functions $\partial^{i}_{x}f$. By solving Thus, solving (3.2) for each $t\in[T]$, we obtain a sequence $w(1),\ldots,w(T)\in\mathbb{R}^{k}$ and consequently, we get a sequence of functions $f(w(1),\cdot),\ldots,f(w(t),\cdot)$. Now, these functions can be evaluated on arbitrarily sets $I_{t}\subset\mathbb{R}$ whose point not necessarily need to be equidistant. Setting $C_{t}=f(w(t),I_{t})+{\epsilon}_{y}\in\mathbb{R}^{|I_{t}|}$, we finally obtain a sequence of process curves $C_{1},\ldots,C_{T}$. A compact overview of the data generation method is shown in Algorithm 1.
Algorithm 1 Generation of process curves
Input: $f:\mathbb{R}^{k}\times\mathbb{R}\to\mathbb{R}$, $x^{i}(1),y^{i}(1),\ldots x^{i}(T),y^{i}(T)\in\mathbb{R}^{n_{i}}$, $i\in[l]$, $\overline{x}\in\mathbb{R}$, $\Delta x\in\mathbb{R}_{>0}$, $m\in\mathbb{N}$
Output: Process curves $C_{1},\ldots,C_{T}$
1:for $t\in[T]$ do
2:     Compute solution $w(t)$ for (3.2) using support points $(x^{1}(t),y^{1}(t)),\ldots,(x^{l}(t),x^{l}(t))$
3:     $I_{t}\leftarrow\{\overline{x}+j\cdot\Delta x+{\epsilon}_{x}:j\in[m]\}$
4:     $C_{t}\leftarrow f(w(t),I_{t})+{\epsilon}_{y}$
5:end for
6:return $C_{1},\ldots,C_{T}$
$x_{j}^{i}(t)$$j$-th support point for $\partial_{x}^{i}f$ with $j\in[n_{i}]$$i$-th derivative of $f$ with $i\in[l]$$t$-th process execution with $t\in[T]$
Fig. 4. Short overview of our notation.
Its left to show how to generate the support points as input for Algorithm 1. One way is to use support points of a real process curve dataset, and using Algorithm 1 to create semi-synthetic copy of it. In a fully synthetic setting, the support points at execution $t\in[T]$, the support points $(x^{i}(t),y^{i}(t))$ can be sampled from a distribution on $\mathbb{R}^{n_{i}}$ respectively, whose statistical properties change over the temporal axis. For instance, $y^{i}(t)\sim\mathcal{N}_{\mu^{i}(t),\sigma}$ with $\mu^{i}:[T]\to\mathbb{R}^{n_{i}}$ encoding the drift behavior over the temporal axis for the support points. Another free parameter of Algorithm 1 is the function $f$ to use. In principle, $f$ can be chosen from any parametrized function set, like B-splines, Gaussian processes [33], neural networks [34], or Kolmogorov-Arnold networks [35]. In Appendix B, we showcase in depth an example where $f$ is a polynomial.
@@
Our theoretic framework extends naturally to multivariate time series data, where each dimension $d\in[c]$ (or signal) has its own function $f_{d}$. If they do not share their latent information $w_{d}(t)$, then Algorithm 1 can be executed for each dimension individually. If they share some latent information, then (3.2) can be extended by summing all support point conditions for all $f_{1},\ldots,f_{c}$.
Remark 3.2 (Profile data).
Our theoretic framework is also capable to generate profile data with drifts holding both, drifts within a profile and drifts over executions. This can be obtained, for instance, by describing how the support points should behave in each profile and for subsequent profiles.